# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v1

@@ 3.2 Busemann points and 1-Lipschitz and Lipschitz functions
The positive part $\delta=\operatorname{RFunk}^{+}=\max(\operatorname{RFunk},0)$ defines a weak metric. See [PT09, Wal08] and [Fun29]. The Thompson metric on the interior of $C$ is defined by
$\operatorname{Th}(x,y)=\max(\operatorname{RFunk}(x,y),\operatorname{RFunk}(y,x)).$
The topology of the Thompson metric coincides with the Euclidean topology on the interior of $C$. We note that the symmetrized metric $\delta_{s}(x,y)=\operatorname{RFunk}^{+}(x,y)+\operatorname{RFunk}^{+}(y,x)$ satisfies $\operatorname{Th}(x,y)\leq\delta_{s}(x,y)\leq 2\operatorname{Th}(x,y)$. By a standard result (e.g. proof of Lemma 4.1 of [AGN16]) a map $f:\operatorname{int}C\to\mathbb{R}$ is non-expansive in the $\operatorname{RFunk}^{+}$ weak metric iff it satisfies:
$x\leq y\implies f(x)\leq f(y),\qquad f(\alpha x)\leq\log\alpha+f(x),\;\text{ for }x,y\in\operatorname{int}{C},\;\alpha\geq 1.$
Maps of this nature arise in game theory, being the coordinates of Shapley operators (dynamic programming operators of zero-sum games) [AGH18]. Observe first that any isometry in the $\operatorname{RFunk}^{+}$ metric is an isometry of Thompson metric. The group of isometries of Thompson metric has been characterized for several cones. If $C=\mathbb{R}_{\geq}^{n}$, any Thompson-isometry is either of the form $x\mapsto T(x)=DPx$ where $D$ is a diagonal matrix with positive diagonal entries and $P$ is a permutation matrix, or of the form $x\mapsto T(x^{-1})$ where $\cdot^{-1}$ denotes the entrywise inverse. Indeed, the map $\log:\mathbb{R}_{>0}^{n}\to\mathbb{R}^{n}$ which takes the entrywise logarithm is an isometry from the open cone $\mathbb{R}_{>0}^{n}$ equipped with Thompson metric to $\mathbb{R}^{n}$ equipped with the sup-norm. Then, this characterization follows from the Banach-Stone theorem. When in addition we assume that $T$ is an isometry with respect to $\operatorname{RFunk}^{+}$, the map $T$ is order preserving, and so the case $x\mapsto T(x^{-1})$ is excluded, so $T(x)=DPx$. Then, it follows from 34 that every isomorphism of the space $\operatorname*{\operatorname{Lip}_{1}}(\mathbb{R}_{>0}^{n},\mathbb{R})$ with respect to the $\operatorname{RFunk}^{+}$ weak metric is of the form
$f\mapsto g(\cdot)+f(DP\cdot).$
where $g:\mathscr{X}\to\mathbb{R}$ is nonexpansive for $\operatorname{RFunk}^{+}$ and $g(x)-g(x^{\prime})+\operatorname{RFunk}^{+}(DPx,DPx^{\prime})=\operatorname{RFunk}^{+}(x,x^{\prime})$. Similarly, if $C=S_{n}^{+}$ is the cone of $n\times n$ real positive semidefinite matrices, then, every Thompson-isometry is of the form $T(x)=axa^{*}$ or $T(x)=ax^{-1}a^{*}$ for some invertible matrix $a$, see [Mol09, Theorem 1]. The second situation is excluded when $T$ is a $\operatorname{RFunk}^{+}$ isometry, hence, by 34 any isomorphism of the space $\operatorname*{\operatorname{Lip}_{1}}(\operatorname{int}S_{n}^{+},\mathbb{R})$ with respect to the $\operatorname{RFunk}^{+}$ weak metric is of the form
$f\mapsto g(\cdot)+f(a\cdot a^{*}).$
where $g:\mathscr{X}\to\mathbb{R}$ is nonexpansive for $\operatorname{RFunk}^{+}$ and, $g(x)-g(x^{\prime})+\operatorname{RFunk}^{+}(axa^{*},ax^{\prime}a^{*})=\operatorname{RFunk}^{+}(x,x^{\prime})$.
As already discussed in 24, inf-stable complete subspaces of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ are sets of non-expansive functions for the $\overline{\mathbb{R}}$-valued weak metric given by $e_{x}$. We can thus reformulate 37 in a more abstract form:
Corollary 41.
Let $\mathcal{G}$ be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ satisfying 1 and such that, for all $x\in\mathscr{X}$, the function $e_{x}$ is finite-valued. Then the set of inf-irreducible elements of $\overline{\mathcal{G}}^{\inf}$ consists of the limits of almost-geodesics up to translations by additive constants, among them the $\{e_{x}+\lambda\}_{x\in\mathscr{X},\,\lambda\in\overline{\mathbb{R}}}$. If for all $x\in\mathscr{X}$, we have $e_{x}\geq 0$ and that $d(x,x^{\prime})=e_{x}(x^{\prime})+e_{x^{\prime}}(x)$ is such that i) $(\mathscr{X},d)$ has compact balls or ii) $e_{x}(x^{\prime})=e_{x^{\prime}}(x)$ and $(\mathscr{X},d)$ is complete, then the functions $\{e_{x}+\lambda\}_{x\in\mathscr{X},\,\lambda\in\overline{\mathbb{R}}}$ are the only inf-irreducible elements of $\overline{\mathcal{G}}^{\inf}$ with maximal Archimedean class.
Proof.
The first part is a reformulation of 36 using that $e_{x}$ is a $\overline{\mathbb{R}}$-valued weak metric. The second is a consequence of 38. ∎
Moving on from 1-Lipschitz to Lipschitz functions, the result below first appeared for order isomorphisms over bounded metric spaces in [SS11, Theorem 1] with a proof in three parts. Adding the assumption of commutation with addition of scalars, we simplified the first part of their proof, about showing the structure of $J$, by considering isomorphisms of lower semicontinuous functions and leveraging Dirac masses. The idea of using Baire’s theorem can be found in [SS11, Lemma 4], here we use seminorms instead of norms which allows us to remove their assumption of boundedness of $\mathscr{X}$ and of $\mathscr{Y}$.
Theorem 42 ((max,+)-isomorphisms of Lipschitz).
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be the set of Lipschitz functions over a complete metric space $(\mathscr{X},d)$ (resp. $(\mathscr{Y},d^{\prime})$). Then every (max,+)-isomorphism $J$ from $\mathscr{F}$ onto $\mathcal{G}$ is of the form
$Jf(x)=g(x)+f(\phi(x))$ (35)
where $g:\mathscr{X}\rightarrow\mathbb{R}$ is a Lipschitz function and $\phi:\mathscr{X}\rightarrow\mathscr{Y}$ is a bi-Lipschitz homeomorphism, i.e. $\phi$ and $\phi^{-1}$ are both Lipschitz.
Proof.
By 15, since Lipschitz functions are dense in lower semicontinuous functions for the topology of pointwise convergence, $J$ has to be of the form given by (25), i.e. $Jf=g+f\circ\phi$ with $\phi$ a homeomorphism and $g$ continuous. Clearly, $g=J(x\mapsto 0)$ is Lipschitz. Define the seminorm $\|\cdot\|_{\text{Lip}(\mathscr{X})}$ over $\text{Lip}(\mathscr{X})$ as $\|f\|_{\text{Lip}(\mathscr{X})}=\sup_{x\neq y}\frac{f(x)-f(y)}{d(x,y)}$. We are going to show that $J$ sends an open ball on a bounded set. By [Kel75, Theorem 34, p200], complete pseudometric spaces are Baire spaces. Seminorms induce pseudometrics, hence $(\text{Lip}(\mathscr{X}),\|\cdot\|_{\text{Lip}(\mathscr{X})})$ is a Baire space. We know that
$\text{Lip}(\mathscr{Y})=\bigcup_{\kappa\in\mathbb{N}^{*}}\{f\in\text{Lip}(\mathscr{Y})\,|\,\|Jf\|_{\text{Lip}(\mathscr{X})}\leq\kappa\}.$
Let us show that these are closed subsets. Fix $\kappa\in\mathbb{N}^{*}$ and take $(f_{n})_{n\in\mathbb{N}}$ with $\|Jf_{n}\|_{\text{Lip}(\mathscr{X})}\leq\kappa$, converging in $(\text{Lip}(\mathscr{Y}),\|\cdot\|_{\text{Lip}(\mathscr{Y})})$ to some $\bar{f}$. Fix $\bar{y}\in\mathscr{Y}$, then $(f_{n}-f_{n}(\bar{y}))_{n}$ converges pointwise to $(\bar{f}-\bar{f}(\bar{y}))$. Since $\phi$ is a homeomorphism, we deduce that $J(f_{n}-f_{n}(\bar{y}))$ also converges pointwise to $J(\bar{f}-\bar{f}(\bar{y}))$ and, since $\|J(f_{n}-f_{n}(\bar{y}))\|_{\text{Lip}(\mathscr{X})}=\|Jf_{n}\|_{\text{Lip}(\mathscr{X})}\leq\kappa$, we obtain that $\|J\bar{f}\|_{\text{Lip}(\mathscr{X})}\leq\kappa$. In other words the sets are closed and by Baire’s theorem, we can fix $\kappa>0$ such that $\{f\in\text{Lip}(\mathscr{Y})\,|\,\|Jf\|_{\text{Lip}(\mathscr{X})}\leq\kappa\}$ has a non-empty interior, and thus contains a closed ball $\mathbb{B}_{\text{Lip}(\mathscr{Y})}(f_{0},\kappa_{0})$ with $\kappa_{0}>0$. By linearity of $J$ w.r.t. $f$, for any $f\in\mathbb{B}_{\text{Lip}(\mathscr{Y})}(0,\kappa_{0})$, we have
$Jf=g+f\circ\phi=g-(g+f_{0}\circ\phi)+g+(f+f_{0})\circ\phi=g-Jf_{0}+J(f+f_{0}).$
Taking the Lipschitz seminorm, this leads to
$J(\mathbb{B}_{\text{Lip}(\mathscr{Y})}(0,\kappa_{0}))\subset\{h\in\text{Lip}(\mathscr{X})\,|\,\|h\|_{\text{Lip}(\mathscr{X})}\leq\|g\|_{\text{Lip}(\mathscr{X})}+\|Jf_{0}\|_{\text{Lip}(\mathscr{X})}+\kappa=:K_{0}\}.$
Since $J$ sends Lipschitz functions on Lipschitz functions, we have that, for all Lipschitz functions $f:\mathscr{Y}\rightarrow\mathbb{R}$ and all $x,x^{\prime}\in\mathscr{X}$
$Jf(x)-Jf(x^{\prime})=g(x)-g(x^{\prime})+f(\phi(x))-f(\phi(x^{\prime}))\leq\|Jf\|_{\text{Lip}(\mathscr{X})}d(x,x^{\prime})$ (36)
In particular, setting $f=\kappa_{0}d^{\prime}(\cdot,\phi(x^{\prime}))$, we get that for all $x,x^{\prime}\in\mathscr{X}$
$-\|g\|_{\text{Lip}(\mathscr{X})}d(x,x^{\prime})+\kappa_{0}d^{\prime}(\phi(x),\phi(x^{\prime}))\leq g(x)-g(x^{\prime})+\kappa_{0}d^{\prime}(\phi(x),\phi(x^{\prime}))\\ \leq\|J(\kappa_{0}d^{\prime}(\cdot,\phi(x^{\prime})))\|_{\text{Lip}(\mathscr{X})}d(x,x^{\prime})\leq K_{0}d(x,x^{\prime}).$ (37)
Hence, $\kappa_{0}d^{\prime}(\phi(x),\phi(x^{\prime}))\leq(\|g\|_{\text{Lip}(\mathscr{X})}+K_{0})d(x,x^{\prime})$, and $\phi$ is Lipschitz. Repeating the proof for $J^{-1}$ shows that $\phi^{-1}$ is also Lipschitz.
@@ 3.3 Continuous affine maps and the lower semicontinuous convex functions
We conclude studying the order isomorphisms over the set of lower semicontinuous convex functions, extending the result of [AAM09] to the generality of locally convex Hausdorff spaces. Our novel ingredient is that we identify in 43 the sup-irreducible points of the space of proper convex lower semicontinuous functions.
@@
$\Rightarrow$: Take $f$ to be a sup-irreducible proper convex l.s.c. function. From [AB06, Theorem 7.6] we know that
@@
denoting by $P^{f}\subset\mathscr{X}^{*}\times\mathbb{R}$ the nonempty set over which the supremum is taken. Fix $x_{0},x_{1}\in\mathscr{X}$, being possibly equal, and $(p_{0},\lambda_{0})\in P^{f}$, $(p_{1},\lambda_{1})\in P^{f}$. Assume that $p_{0}\neq p_{1}$, then by Hahn-Banach’s theorem applied to the locally convex space $\mathscr{X}^{*}$ equipped with the weak-* topology, we can take a weak-* closed halfspace $H^{+}$ containing $p_{0}$ but not $p_{1}$ and denote by $H^{-}$ the closure of its complement. This means that $H^{-}=\{p\in\mathscr{X}^{*}\,|\,\langle p,x_{H}\rangle\leq\alpha_{H}\}$ for some $x_{H}\in\mathscr{X}\backslash\{0\}$ and $\alpha_{H}\in\mathbb{R}$, w.l.o.g  up to a translation in $p$, we can assume $\alpha_{H}=0$. Similarly, up to a translation in $x$, we can assume $f(0)\in\mathbb{R}$.
@@
Let $g(x):=\sup_{(p,\lambda)\in(H^{+}\times\mathbb{R})\cap P^{f}}\langle p,x\rangle+\lambda$ and $h(x):=\sup_{(p,\lambda)\in(H^{-}\times\mathbb{R})\cap P^{f}}\langle p,x\rangle+\lambda$. Then clearly $f=\sup(g,h)$ but $f\neq g$ and $f\neq h$. Indeed, since $p_{0}\notin H^{-}$, $\langle p_{0},x_{H}\rangle>0$. Consequently we can fix $\beta\in\mathbb{R}_{+}^{*}$ such that $\beta\langle p_{0},x_{H}\rangle\geq f(0)-\lambda_{0}$. For all $(p,\lambda)\in(H^{-}\times\mathbb{R})\cap P^{f}$, we deduce from (38) that $f(0)\geq\lambda$ and we have that $\langle p_{0},x_{H}\rangle>0\geq\langle p,x_{H}\rangle$, whence $\langle p_{0},\beta x_{H}\rangle+\lambda_{0}\geq f(0)\geq\langle p,\beta x_{H}\rangle+\lambda$ since $\langle p,\beta x_{H}\rangle\leq 0$. Since $(p_{0},\lambda_{0})\in P^{f}$, taking the supremum in $(H^{-}\times\mathbb{R})\cap P^{f}$ and fixing $y_{0}=\beta x_{h}$ shows that $h(y_{0})\leq\langle p_{0},y_{0}\rangle+\lambda_{0}<f(y_{0})$ so $h\neq f$. The same argument applies to $g$. Hence, $p_{0}=p_{1}$, and
$f(x)=\sup_{\lambda\in\mathbb{R}}\{\langle p_{0},x\rangle+\lambda\,|\,\forall\,y\in\mathscr{X},\,\langle p_{0},y\rangle+\lambda<f(y)\}=\langle p_{0},x\rangle-\sup_{y}[\langle p_{0},y\rangle-f(y)],$
so $f$ is a continuous affine function. ∎
@@
The following theorem has previously been shown for $\mathscr{X}=\mathbb{R}^{d}$ in [AAM09, Corollary 8]. Its generalization to Banach spaces appeared in [IRS15, Theorem 1]. All proofs rely eventually on a key result of affine geometry: in dimension greater than 2 only affine maps preserve segments. It should be highlighted that it is already stated in [AAM09, p670], before their Remark 9, that the characterization of anti-involutions given in 45 could be obtained from that of order isomorphisms as in 44. Artstein-Avidan and Milman tackle first anti-involutions and then the isomorphisms. The difference between the two is that isomorphisms send relatively-inf-irreducible elements (here the indicator functions) onto themselves as shown in 13, while anti-involutions send them onto the sup-irreducible elements (here the continuous affine functions). This explains the main discrepancies of our proof with that of [AAM09, Theorem 5] and its similarity with that of [IRS15, Theorem 1]. To the difference of [IRS15], an argument based on weak-* continuity enables us to avoid all together considering the bidual in the proof. Aside from this result, our proof method is different because we emphasize the role of the sup-irreducible elements to obtain the characterization. Concerning anti-involutions, [CL21, Theorem 1.3] shows that reflexivity and a linear isomorphism between $\mathscr{X}$ and $\mathscr{X}^{*}$ are necessary and sufficient to have the Fenchel transform as anti-involution, as otherwise the bi-Fenchel is defined on the larger set $\mathscr{X}^{**}$.
@@
Let $\mathcal{G}$ be the space of proper convex lower semicontinuous functions over a locally convex Hausdorff topological vector space $\mathscr{X}$ of dimension larger than two, then the isomorphisms over $\mathcal{G}$ are affine, i.e. there exists a weakly-weakly continuous linear map $A:\mathscr{X}\to\mathscr{X}$ invertible with weakly-weakly continuous inverse, and some constants $c\in\mathscr{X}$, $b\in\mathscr{X}^{*}$, $d,\delta\in\mathbb{R}$ with $d>0$ such that, for all $f\in\mathcal{G}$, we have
$Jf(x)=\langle b,x\rangle+\delta+d\cdot f(Ax+c).$ (39)
@@
Using 13 and 43, we have that $J$ maps continuous affine functions onto continuous affine functions since they are the sup-irreducible points in this case. Let $G:\mathscr{X}^{*}\times\mathbb{R}\rightarrow\mathscr{X}^{*}\times\mathbb{R}$ be the bijective function that maps the coefficients $(p,\lambda)$ of an affine function $f$ onto the coefficients $(p^{\prime},\lambda^{\prime})$ of $Jf$. Let us write $G(p,\lambda)=(G_{1}(p,\lambda),G_{2}(p,\lambda))$
Take any two affine functions $f_{0}(\cdot):=\langle p_{0},\cdot\rangle+\lambda_{0}$ and $f_{1}(\cdot):=\langle p_{1},\cdot\rangle+\lambda_{1}$ for some $p_{0},p_{1}\in\mathscr{X}^{*}$, $\lambda_{0},\lambda_{1}\in\mathbb{R}$. Fix any $\lambda^{\prime}_{0}\in\mathbb{R}$, w.l.o.g. assume that $\lambda^{\prime}_{0}>\lambda_{0}$. Since $J$ preserves the order, $J(f_{0})\leq J(f_{0}+\lambda^{\prime}_{0}-\lambda_{0})$, both functions being affine, they must have the same $p$-coefficient, which thus does not depend on $\lambda$. In other words, $G_{1}(\cdot,\lambda)=G_{1}(\cdot,0)$.
@@
with equality at some point $x_{0}$. Consequently, since $J$ commutes with the max operation, we have that $Jf_{\alpha}\leq\max(Jf_{0},Jf_{1})$ with all functions being affine. Let us show that there exists a point $x^{\prime}_{0}$ where there is equality.
If $Jf_{\alpha}<\max(Jf_{0},Jf_{1})$, then, $Jf_{\alpha}$ being continuous affine and the r.h.s. being a max of continuous affine functions, Hahn-Banach’s strict separation entails there exists $\epsilon>0$ such that

### v2

@@ 3.2 Busemann points and 1-Lipschitz functions
It is a weak metric, see [PT09, Wal08] and [Fun29]. The Hilbert metric on the interior of $C$ is defined by
$\operatorname{Hil}(x,y)=\operatorname{RFunk}(x,y)+\operatorname{RFunk}(y,x).$
It will be convenient to fix a linear form $p$ in the interior of the dual cone of $C$, and to consider the cross-section $C_{p}=\{x\in\operatorname{int}{C}\mid p(x)=1\}$. Then, $(x,y)\mapsto\operatorname{Hil}(x,y)$ is a bona fide metric on $C_{p}$. Observe that all the Hilbert balls of $C_{p}$ are compact, so that 34 can be applied. The group of isometries of $(C_{p},\operatorname{Hil})$ has been characterized in [LW11, Theorem 1.2] when $C=\mathbb{R}_{\geq}^{n}$ is the standard orthant, so that $C_{p}$ is the simplex. Let $\jmath$ denote the self-map of $(\mathbb{R}_{>0})^{n}$ such that $(\jmath(x))_{i}=x_{i}^{-1}$ for all $i\in[n]$, and $\rho$ denote the map from $\mathbb{R}_{>0}^{n}$ to $C_{p}$ such that $\rho(x)=x/p(x)$. Then, every isometry of $(C_{p},\operatorname{Hil})$ is either of the form $\phi(x)=\rho(DPx)$ where $D$ is a diagonal matrix with positive diagonal entries and $P$ is a permutation matrix, or of the form $\phi(x)=\rho(\jmath(DPx))$. Then, 34 shows that every (max,+)-isomorphism of the space of nonexpansive maps defined on $(C_{p},\operatorname{Hil})$ is of the form (30) where the function $g$ is constant. 34 also allows to characterize the set of (max,+)-isomorphisms of the space of real nonexpansive maps defined on $(C_{p},\operatorname{RFunk})$. When $n\geq 3$, considering $\delta=\delta^{\prime}=\operatorname{RFunk}$ in (30), it is straightforward to check that $g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))=\delta(x,x^{\prime})$ holds for all $x,x^{\prime}\in C_{p}$ iff $g$ is constant and $\phi$ is of the form $\phi(x)=\rho(DPx)$ (the case in which $\phi=\rho\circ\jmath(DPx)$ is excluded owing to the “unilateral” character of the Funk metric).
@@ 3.3 Continuous affine maps and the lower semicontinuous convex functions
We conclude studying the order isomorphisms over the set of lower semicontinuous convex functions, extending the result of [AAM09] to the generality of locally convex Hausdorff spaces. Our novel ingredient is that we identify in 41 the sup-irreducible points of the space of proper convex lower semicontinuous functions.
@@
$\Rightarrow$: Take $f$ to be a sup-irreducible proper convex l.s.c. function. Fixing $\bar{x}\in\operatorname*{Dom}(f)$, we can assume w.l.o.g. $f(0)\in\mathbb{R}$ by replacing $f$ by $f(\cdot-\bar{x})$ which is also sup-irreducible. From [AB06, Theorem 7.6] we know that
@@
denoting by $P^{f}\subset\mathscr{X}^{*}\times\mathbb{R}$ the nonempty set over which the supremum is taken. Fix $x_{0},x_{1}\in\mathscr{X}$, being possibly equal, and $(p_{0},\lambda_{0})\in P^{f}$, $(p_{1},\lambda_{1})\in P^{f}$. Assume that $p_{0}\neq p_{1}$, then by Hahn-Banach’s theorem applied to the locally convex space $\mathscr{X}^{*}$ equipped with the weak-* topology, we can take a weak-* closed halfspace $H^{+}$ containing $p_{0}$ but not $p_{1}$ and denote by $H^{-}$ the closure of its complement. This means that $H^{-}=\{p\in\mathscr{X}^{*}\,|\,\langle p,x_{H}\rangle\leq\alpha_{H}\}$ for some $x_{H}\in\mathscr{X}\backslash\{0\}$ and $\alpha_{H}\in\mathbb{R}$.
Let us get rid of $\alpha_{H}$ first. Set $\tilde{f}=f-\langle\bar{p},\cdot\rangle$ where $\bar{p}\in\mathscr{X}^{*}$ is such that $\langle\bar{p},x_{H}\rangle=\alpha_{H}$. Then $\tilde{f}$ is sup-irreducible and $P^{\tilde{f}}=P^{f}-(\bar{p},0)$. So replacing $p_{0}$ by $p_{0}-\bar{p}$ and $p_{1}$ by $p_{1}-\bar{p}$, both are still separated, but this time by $H^{-}=\{p\in\mathscr{X}^{*}\,|\,\langle p,x_{H}\rangle\leq 0\}$.
@@
Let $g(x):=\sup_{(p,\lambda)\in(H^{+}\times\mathbb{R})\cap P^{\tilde{f}}}\langle p,x\rangle+\lambda$ and $h(x):=\sup_{(p,\lambda)\in(H^{-}\times\mathbb{R})\cap P^{\tilde{f}}}\langle p,x\rangle+\lambda$. Then clearly $\tilde{f}=\sup(g,h)$ but $\tilde{f}\neq g$ and $\tilde{f}\neq h$. Indeed, since $p_{0}\notin H^{-}$, $\langle p_{0},x_{H}\rangle>0$. Consequently we can fix $\beta\in\mathbb{R}_{+}^{*}$ such that $\beta\langle p_{0},x_{H}\rangle\geq\tilde{f}(0)-\lambda_{0}$. For all $(p,\lambda)\in(H^{-}\times\mathbb{R})\cap P^{\tilde{f}}$, we deduce from (36) that $\tilde{f}(0)\geq\lambda$ and we have that $\langle p_{0},x_{H}\rangle>0\geq\langle p,x_{H}\rangle$, whence $\langle p_{0},\beta x_{H}\rangle+\lambda_{0}\geq\tilde{f}(0)\geq\langle p,\beta x_{H}\rangle+\lambda$ since $\langle p,\beta x_{H}\rangle\leq 0$. Since $(p_{0},\lambda_{0})\in P^{\tilde{f}}$, taking the supremum in $(H^{-}\times\mathbb{R})\cap P^{\tilde{f}}$ and fixing $y_{0}=\beta x_{h}$ shows that $h(y_{0})\leq\langle p_{0},y_{0}\rangle+\lambda_{0}<\tilde{f}(y_{0})$ so $h\neq\tilde{f}$. The same argument applies to $g$. Hence, $p_{0}=p_{1}$, and
$\tilde{f}(x)=\sup_{\lambda\in\mathbb{R}}\{\langle p_{0},x\rangle+\lambda\,|\,\forall\,y\in\mathscr{X},\,\langle p_{0},y\rangle+\lambda<f(y)\}=\langle p_{0},x\rangle-\sup_{y}[\langle p_{0},y\rangle-f(y)].$
Hence $\tilde{f}$ is a continuous affine function, and so is $f=\tilde{f}+\langle\bar{p},\cdot\rangle$. ∎
@@
The following theorem has previously been shown for $\mathscr{X}=\mathbb{R}^{d}$ in [AAM09, Corollary 8]. Its generalization to Banach spaces appeared in [IRS15, Theorem 1]. All proofs rely eventually on a key result of affine geometry: in dimension greater than 2 only affine maps preserve segments. It should be highlighted that it is already stated in [AAM09, p670], before their Remark 9, that the characterization of anti-involutions given in 43 could be obtained from that of order isomorphisms as in 42. Artstein-Avidan and Milman tackle first anti-involutions and then the isomorphisms. As shown in 13, the difference between the two is that isomorphisms send relatively-inf-irreducible elements (here the indicator functions) onto themselves, while anti-involutions send them onto the sup-irreducible elements (here the continuous affine functions). This explains the main discrepancies of our proof with that of [AAM09, Theorem 5] and its similarity with that of [IRS15, Theorem 1]. To the difference of [IRS15], an argument based on weak-* continuity enables us to avoid all together considering the bidual in the proof. Aside from this result, our proof method is different because we emphasize the role of the sup-irreducible elements to obtain the characterization. Concerning anti-involutions, [CL21, Theorem 1.3] shows that reflexivity and a linear isomorphism between $\mathscr{X}$ and $\mathscr{X}^{*}$ are necessary and sufficient to have the Fenchel transform as anti-involution, as otherwise the bi-Fenchel is defined on the larger set $\mathscr{X}^{**}$.
@@
Let $\mathcal{G}$ be the space of proper convex lower semicontinuous functions over a locally convex Hausdorff topological vector space $\mathscr{X}$, then the order isomorphisms over $\mathcal{G}$ are affine, i.e. there exists a linear map $A:\mathscr{X}\to\mathscr{X}$ with w*-w* continuous adjoint, which has an inverse with w*-w* continuous adjoint, and some constants $c\in\mathscr{X}$, $b\in\mathscr{X}^{*}$, $d,\delta\in\mathbb{R}$ with $d>0$ such that, for all $f\in\mathcal{G}$, we have
$Jf(x)=\langle b,x\rangle+\delta+d\cdot f(Ax+c),$ (37)
and these $J$ are the only order isomorphisms over $\mathcal{G}$.
As shown in the proof, the existence of a w*-w* continuous adjoint implies both that the linear map is w-w continuous and has closed graph. If $\mathscr{X}$ is Fréchet, $A$ is then continuous with continuous inverse.
@@
Using 13 and 41, we have that $J$ maps continuous affine functions onto continuous affine functions since they are the sup-irreducible points in this case. Let $G:\mathscr{X}^{*}\times\mathbb{R}\rightarrow\mathscr{X}^{*}\times\mathbb{R}$ be the bijective function that maps the coefficients $(p,\lambda)$ of an affine function $f$ onto the coefficients $(p^{\prime},\lambda^{\prime})$ of $Jf$. Let us write $G(p,\lambda)=(G_{1}(p,\lambda),G_{2}(p,\lambda))$
Take any two affine functions $f_{0}(\cdot):=\langle p_{0},\cdot\rangle+\lambda_{0}$ and $f_{1}(\cdot):=\langle p_{1},\cdot\rangle+\lambda_{1}$ for some $p_{0}\neq p_{1}\in\mathscr{X}^{*}$, $\lambda_{0},\lambda_{1}\in\mathbb{R}$. Fix any $\lambda^{\prime}_{0}\in\mathbb{R}$, w.l.o.g. assume that $\lambda^{\prime}_{0}>\lambda_{0}$. Since $J$ preserves the order, $Jf_{0}\leq J(f_{0}+\lambda^{\prime}_{0}-\lambda_{0})$, both functions being affine, they must have the same $p$-coefficient, which thus does not depend on $\lambda$. In other words, $G_{1}(\cdot,\lambda)=G_{1}(\cdot,0)$.
@@
with equality at some point $x_{0}$. Consequently, since $J$ commutes with the max operation by 7, we have that $Jf_{\alpha}\leq\max(Jf_{0},Jf_{1})$ with all functions being affine. Let us show that there exists a point $x^{\prime}_{0}$ where there is equality.
If $Jf_{\alpha}<\max(Jf_{0},Jf_{1})$, let us show there exists $\epsilon>0$ such that